# IDT: A Physically Grounded Transformer for Feed-Forward Multi-View Intrinsic Decomposition

### v1

@@ Abstract
Intrinsic image decomposition is fundamental for visual understanding, as RGB images entangle material properties, illumination, and view-dependent effects. Recent diffusion-based methods have achieved strong results for single-view intrinsic decomposition; however, extending these approaches to multi-view settings remains challenging, often leading to severe view inconsistency. We propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for multi-view intrinsic image decomposition. By leveraging transformer-based attention to jointly reason over multiple input images, IDT produces view-consistent intrinsic factors in a single forward pass, without iterative generative sampling. IDT adopts a physically grounded image formation model that explicitly decomposes images into diffuse reflectance, diffuse shading, and specular shading. This structured factorization separates Lambertian and non-Lambertian light transport, enabling interpretable and controllable decomposition of material and illumination effects across views. Experiments on both synthetic and real-world datasets demonstrate that IDT achieves cleaner diffuse reflectance, more coherent diffuse shading, and better-isolated specular components, while substantially improving multi-view consistency compared to prior intrinsic decomposition methods.
@@ 1 Introduction
Intrinsic image decomposition aims to separate an image into its underlying material and illumination components, providing a more interpretable representation than raw RGB images. Such decomposition is fundamental for a wide range of visual tasks, including relighting, material editing, 3D reconstruction, and world modeling, where entangled appearance factors often limit generalization and controllability [Bell2014Intrinsic]. Despite decades of research, intrinsic decomposition remains challenging due to the complex interaction between geometry, illumination, and view-dependent effects.
@@
Recent advances in large-scale learning have significantly improved intrinsic decomposition from a single image. Learning-based methods, including convolutional and transformer-based architectures, have demonstrated strong performance under single-view settings [Baslamisli2018Intrinsic]. Beyond pure RGB inputs, several works have explored incorporating additional geometric or modality cues (e.g., depth or surface normals) to reduce intrinsic ambiguities, following the broader RGB-X paradigm. More recently, diffusion-based models have further advanced single-view intrinsic decomposition by leveraging powerful generative priors and iterative refinement. However, these approaches typically operate on each image independently and are primarily designed for per-view appearance generation, without explicitly enforcing consistency across views.
Extending intrinsic decomposition to multi-view settings introduces a fundamental challenge: view consistency. Independently decomposed views often exhibit inconsistencies in material appearance and illumination, especially in the presence of view-dependent effects such as specular highlights. Such inconsistencies severely limit the applicability of intrinsic decomposition in multi-view reconstruction and downstream 3D scene understanding. Moreover, iterative generative formulations are not naturally suited for enforcing cross-view coherence, as they lack a mechanism for joint reasoning across multiple observations.
@@
In this work, we take a different perspective. We build upon recent progress in feed-forward geometric reasoning with transformers, where multi-view images are jointly processed to infer consistent scene geometry. Transformer-based architectures such as VGGT demonstrate that attention mechanisms can effectively aggregate information across views and produce coherent geometric representations in a single forward pass [Wang2025VGGT]. Motivated by this property, we leverage transformer attention to perform feed-forward, multi-view intrinsic decomposition, jointly reasoning over multiple input images without iterative sampling.
@@
Beyond multi-view consistency, we emphasize the importance of physically grounded intrinsic decomposition, particularly in indoor environments. Rather than treating intrinsic components as abstract latent factors or unstructured residuals, we explicitly model image formation as the combination of diffuse reflectance (albedo), diffuse shading, and specular shading. This formulation is consistent with physically motivated rendering decompositions adopted in large-scale indoor datasets such as Hypersim [Roberts2021Hypersim], and enforces a clear separation between Lambertian and non-Lambertian light transport. As a result, the resulting intrinsic representations are more interpretable and controllable across different viewpoints.
@@
Based on these insights, we propose Intrinsic Decomposition Transformer (IDT), a feed-forward multi-view framework that unifies transformer-based geometric reasoning with a physically grounded intrinsic formulation. IDT jointly infers view-consistent albedo, diffuse shading, and specular shading from multiple input images, enabling robust intrinsic decomposition in challenging multi-view indoor scenarios.
Our contributions are summarized as follows:
@@
  • We introduce a feed-forward transformer framework for view-consistent multi-view intrinsic image decomposition, inspired by recent advances in multi-view geometric reasoning.
@@
  • We propose a physically grounded intrinsic formulation that explicitly separates diffuse reflectance, diffuse shading, and specular shading, leading to interpretable intrinsic factors.
@@
  • We demonstrate improved multi-view consistency and decomposition quality on both synthetic and real-world indoor datasets compared to prior intrinsic decomposition methods.
@@ 3.2 Feed-Forward Multi-View Intrinsic Decomposition
where albedo $\mathbf{A}$ and illumination $\mathbf{L}$ are shared across all views, while shading components remain view-dependent. This feed-forward formulation enforces cross-view consistency at the representation level and avoids iterative generative inference commonly used in prior intrinsic and inverse rendering approaches, as shown in LABEL:1111.
@@ 3.4 Appearance Adapters for Intrinsic Prediction
Appearance-related representations produced by the multi-view transformer are expressive but highly redundant. In multi-view intrinsic decomposition, different intrinsic factors rely on distinct and partially conflicting cues: diffuse reflectance (albedo) should emphasize view-invariant material properties, whereas diffuse and specular shading are strongly influenced by view-dependent effects and global illumination. Directly feeding the shared token set into all prediction heads often leads to entanglement between material and illumination cues.
To enable selective and factor-specific reasoning, we introduce appearance adapters that route information from the shared token set to each intrinsic head. For each intrinsic factor $k\in\{\text{alb},\text{diff},\text{spec}\}$, an adapter extracts a compact, task-specific representation:
@@
where $\mathbf{Z}$ denotes the aggregated multi-view tokens from the encoder.
@@
Each adapter is implemented as a lightweight cross-attention block. Crucially, the queries are not learned parameters but are derived from the encoder outputs: we construct queries by pooling scene-level tokens (e.g., camera and register tokens) across views, while keys and values are formed from patch tokens across all views. This design allows scene-level representations to attend selectively to multi-view appearance cues, producing a compact scene-conditioned context for each intrinsic factor. By conditioning attention on scene tokens rather than learned slots, the adapter preserves geometric and cross-view coherence while remaining fully feed-forward.
@@ 4 Experiments
Figure 4: Qualitative multi-view intrinsic decomposition results on synthetic and real-world scenes. Given three input images from different viewpoints, IDT performs feed-forward joint inference to produce view-consistent diffuse reflectance, diffuse shading, specular shading, and surface normals. The results highlight effective separation of Lambertian and non-Lambertian effects across views.
Table 1: Quantitative results on the Hypersim dataset. We evaluate intrinsic decomposition accuracy, multi-view consistency, and reconstruction quality. Single-view methods are applied independently to each view. Lower is better for MAE and consistency metrics, while higher is better for PSNR and SSIM.
Method Albedo Diffuse Shading Specular Shading Reconstruction
MAE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ MAE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
IntrinsicNet [Baslamisli2018Intrinsic] (single-view)
IIW-CNN [Bell2014Intrinsic] (single-view)
VGGT (per-view) [Wang2025VGGT]
IDT (Ours)
Table 2: Multi-view consistency on Hypersim. Consistency is measured as the average $\ell_{1}$ difference between intrinsic predictions warped to a reference view. Lower is better.
Method A-Cons. $\downarrow$ S-Cons. $\downarrow$
IntrinsicNet [Baslamisli2018Intrinsic] (SV)
VGGT (per-view) [Wang2025VGGT]
IDT (Ours)
We evaluate the proposed Intrinsic Decomposition Transformer (IDT) on both synthetic and real-world datasets to assess intrinsic decomposition accuracy, multi-view consistency, and generalization. All experiments are conducted under the physically grounded image formation model introduced in Sec. 3, with a particular focus on view-consistent decomposition of material and illumination factors.
@@ 4.1 Datasets
Hypersim & InteriorVerse.
We evaluate our method on both synthetic and real-world indoor datasets. Hypersim [Roberts2021Hypersim] provides physically based renderings with full intrinsic annotations, including diffuse reflectance, diffuse illumination, and non-diffuse residuals, and is used as our primary benchmark for quantitative evaluation and ablations under controlled supervision. InteriorVerse [zhu2022learning] consists of real-world or photorealistic indoor scenes with diverse geometry, materials, and lighting, serving as a challenging testbed to assess generalization and robustness in realistic settings where full intrinsic ground truth is unavailable. For both datasets, we construct multi-view samples by grouping $V$ images observing the same static scene from nearby viewpoints, unless otherwise specified.
@@ 4.2 Evaluation Metrics
We evaluate intrinsic decomposition along three complementary dimensions.
Intrinsic decomposition accuracy.
On Hypersim, where ground-truth intrinsic factors are available, we evaluate diffuse reflectance (albedo), diffuse shading, and specular shading using standard image regression metrics, including mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity (SSIM). For shading-related factors, metrics are additionally computed in the logarithmic domain to account for their high dynamic range.
@@
To quantify cross-view consistency, we measure the agreement of predicted intrinsic factors across different views of the same scene. Specifically, predicted albedo and diffuse shading maps from each view are warped to a reference view using available camera geometry, and the average $\ell_{1}$ difference is computed. Lower values indicate better cross-view consistency.
@@
We also evaluate reconstruction quality by recomposing images using the predicted intrinsic factors according to Eq. (2). Reconstruction quality is measured using PSNR and SSIM on both Hypersim and InteriorVerse. This metric is applicable even when ground-truth intrinsic factors are unavailable and reflects how well the predicted decomposition explains the input images.
@@ 4.3 Baselines
We compare IDT against representative intrinsic decomposition baselines covering both single-view and feed-forward multi-view settings.
Single-view intrinsic decomposition.
We include classic and learning-based single-image intrinsic decomposition methods, which are applied independently to each view: Intrinsic Images in the Wild (IIW-CNN) [Bell2014Intrinsic] and IntrinsicNet [Baslamisli2018Intrinsic]. Although these methods are not designed to enforce multi-view consistency, they provide reference points for per-image decomposition quality.
Diffusion-based intrinsic decomposition.
We also compare against recent diffusion-based intrinsic decomposition approaches, which generate intrinsic factors from a single image using iterative generative sampling. These methods are applied independently to each view and do not explicitly enforce cross-view consistency.
Multi-view feed-forward baseline.
To isolate the effect of joint multi-view reasoning, we construct a strong feed-forward baseline based on VGGT [Wang2025VGGT]. This baseline uses the same backbone, prediction heads, training data, and loss functions as IDT, but processes each view independently without joint attention across views. As a result, any performance difference can be attributed to the proposed joint multi-view inference mechanism rather than network capacity or supervision.
4.4 Implementation Details
IDT is implemented using a transformer-based multi-view encoder with intrinsic adapters and lightweight prediction heads. All models are trained in a fully feed-forward manner using the Adam optimizer, with a fixed learning rate and identical training schedules across methods to ensure fair comparison. Each training sample consists of $V$ views randomly sampled from the same scene, enabling joint multi-view inference during training.
To stabilize training and encourage effective disentanglement of intrinsic factors, we adopt a two-stage training strategy. In the first stage, we freeze the multi-view aggregator and train only the intrinsic adapters and prediction heads for 50 epochs, allowing the model to learn task-specific intrinsic mappings on top of a fixed multi-view representation. In the second stage, we unfreeze the aggregator and jointly fine-tune the entire network for an additional 100 epochs, enabling end-to-end refinement of multi-view representations and intrinsic predictions. Unless otherwise specified, all reported results are obtained from models trained with this two-stage schedule.
Training is performed on 8 NVIDIA H100 GPUs with 80 GB memory per GPU using distributed data-parallel training. All models are trained to convergence under the same computational budget. Unless otherwise specified, batch size, learning rate, optimizer parameters, and architectural hyperparameters are kept consistent across experiments.
During training, we supervise diffuse reflectance (albedo), diffuse shading, and specular shading using the loss functions described in Sec. 3.5. Illumination supervision is applied only when lighting information is available; otherwise, illumination representations are learned implicitly through shading prediction and image reconstruction under the proposed image formation model. Detailed network architectures, training schedules, and hyperparameter settings are provided in the supplementary material for reproducibility.

### v2

@@ Abstract
Intrinsic image decomposition is fundamental for visual understanding, as RGB images entangle material properties, illumination, and view-dependent effects. Recent diffusion-based methods have achieved strong results for single-view intrinsic decomposition; however, extending these approaches to multi-view settings remains challenging, often leading to severe view inconsistency. We propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for multi-view intrinsic image decomposition. By leveraging transformer-based attention to jointly reason over multiple input images, IDT produces view-consistent intrinsic factors in a single forward pass, without iterative generative sampling. IDT adopts a physically grounded image formation model that explicitly decomposes images into diffuse reflectance, diffuse shading, and specular shading. This structured factorization separates Lambertian and non-Lambertian light transport, enabling interpretable and controllable decomposition of material and illumination effects across views. Experiments on both synthetic and real-world datasets demonstrate that IDT achieves cleaner diffuse reflectance, more coherent diffuse shading, and better-isolated specular components, while substantially improving multi-view consistency compared to prior intrinsic decomposition methods. Code is available at https://github.com/dukang/IDT.
@@ 1 Introduction
Intrinsic image decomposition seeks to factor an image into its underlying material and illumination components, offering a more interpretable and controllable representation than raw RGB appearance. Such decomposition is a long-standing goal in computer vision and graphics, and serves as a critical foundation for applications including relighting, material editing, multi-view reconstruction, and world modeling, where entangled appearance factors often hinder generalization and physical reasoning [Bell2014Intrinsic]. Despite decades of research, intrinsic decomposition remains fundamentally challenging due to the intertwined effects of geometry, illumination, and view-dependent appearance.
@@
Recent years have witnessed substantial progress in intrinsic decomposition from a single image. Learning-based approaches, ranging from convolutional models to transformer architectures, have significantly improved per-view decomposition quality under supervised or weakly supervised settings [Baslamisli2018Intrinsic]. Beyond pure RGB inputs, several methods further incorporate geometric cues such as depth or surface normals, following the broader RGB-X paradigm, to alleviate intrinsic ambiguities. More recently, diffusion-based models have demonstrated impressive single-view results by leveraging powerful generative priors and iterative refinement. However, a common limitation of these approaches is that they operate on each image independently, implicitly treating intrinsic decomposition as a per-view appearance generation problem.
In contrast, many downstream tasks demand multi-view intrinsic consistency. When multiple images of the same scene are available, independently decomposed intrinsic factors often vary across viewpoints, leading to inconsistent material appearance and unstable illumination estimates. This issue is particularly severe in indoor environments with complex lighting and view-dependent effects such as specular highlights. Such inconsistencies fundamentally limit the use of intrinsic decomposition in multi-view reconstruction, scene editing, and physically grounded world modeling. Moreover, existing iterative or generative formulations lack an explicit mechanism for jointly reasoning across multiple views, making cross-view coherence difficult to enforce.
@@
In this work, we argue that intrinsic decomposition should be revisited from a multi-view, feed-forward perspective. Our approach builds on recent advances in transformer-based geometric reasoning, where multiple views are jointly processed to infer coherent scene representations in a single forward pass. In particular, models such as VGGT demonstrate that attention mechanisms can effectively aggregate information across views and produce consistent geometric predictions without iterative optimization [Wang2025VGGT]. This observation motivates us to extend feed-forward multi-view reasoning beyond geometry, and into the intrinsic decomposition problem.
@@
Beyond consistency, we further emphasize the importance of physically grounded intrinsic representations. Rather than modeling intrinsic components as abstract latent variables or unstructured residuals, we explicitly adopt a physically motivated image formation model that decomposes appearance into diffuse reflectance (albedo), diffuse shading, and specular shading. This formulation aligns with physically based rendering decompositions used in large-scale indoor datasets such as Hypersim [Roberts2021Hypersim], and enforces a principled separation between Lambertian and non-Lambertian effects. As a result, the predicted intrinsic factors are more interpretable, stable across viewpoints, and suitable for downstream physical reasoning.
@@
Based on these insights, we propose Intrinsic Decomposition Transformer (IDT), a feed-forward framework for multi-view intrinsic image decomposition. IDT jointly processes multiple images of a scene using transformer attention, enabling consistent inference of albedo, diffuse shading, and specular shading across views in a single forward pass. By unifying multi-view geometric reasoning with a physically grounded intrinsic formulation, IDT provides a practical and effective solution for intrinsic decomposition in realistic indoor environments.
Our contributions can be summarized as follows:
@@
  • We introduce a feed-forward transformer framework for view-consistent multi-view intrinsic image decomposition, extending recent advances in multi-view geometric reasoning to intrinsic factorization.
@@
  • We propose a physically grounded intrinsic formulation that explicitly separates diffuse reflectance, diffuse shading, and specular shading, leading to interpretable and stable intrinsic representations.
@@
  • We demonstrate improved single-view accuracy and substantially enhanced multi-view consistency on both synthetic and real-world indoor datasets compared to prior intrinsic decomposition methods.
@@ 3.2 Feed-Forward Multi-View Intrinsic Decomposition
where albedo $\mathbf{A}$ and illumination $\mathbf{L}$ are shared across all views, while shading components remain view-dependent. This feed-forward formulation enforces cross-view consistency at the representation level and avoids iterative generative inference commonly used in prior intrinsic and inverse rendering approaches, as shown in Figure 3.
@@ 3.4 Appearance Adapters for Intrinsic Prediction
Appearance-related representations produced by the multi-view transformer are expressive but highly redundant. In multi-view intrinsic decomposition, different intrinsic factors rely on distinct and partially conflicting cues: diffuse reflectance (albedo) should emphasize view-invariant material properties, whereas diffuse and specular shading are strongly influenced by view-dependent effects, lighting direction, and global illumination. Directly feeding the same shared token set into all prediction heads therefore tends to entangle material and illumination cues, often resulting in color leakage in albedo or unstable shading predictions across views.
To enable selective and factor-specific reasoning, we introduce appearance adapters that explicitly route information from the shared multi-view token set to each intrinsic prediction head. Rather than enforcing a single shared representation, the adapters act as lightweight interfaces that extract task-relevant appearance information while suppressing irrelevant or conflicting cues. For each intrinsic factor $k\in\{\text{alb},\text{diff},\text{spec}\}$, an adapter produces a compact, task-specific representation:
@@
where $\mathbf{Z}$ denotes the aggregated multi-view tokens output by the encoder.
@@
Each appearance adapter $\mathcal{A}_{k}$ is implemented as a lightweight cross-attention block. Crucially, the queries are not learnable parameters but are derived directly from the encoder outputs. Specifically, we construct queries by pooling scene-level tokens (e.g., camera tokens and register tokens) across views, which encode global scene geometry and camera configuration. Keys and values are formed from patch-level tokens across all views, representing dense multi-view appearance observations. This asymmetric design allows scene-level representations to selectively attend to relevant appearance cues distributed over multiple views.
By conditioning attention on scene tokens rather than learned slots, the adapter remains fully feed-forward and avoids introducing additional latent variables or iterative optimization. Moreover, because the queries encode scene-level geometry and camera information, the resulting attended features are naturally aligned across views, yielding a scene-conditioned context that is consistent under viewpoint changes. This design encourages the albedo adapter to focus on view-invariant material properties, while allowing the shading adapters to aggregate view-dependent illumination cues in a controlled manner.
Factor-specific decoupling.
Although all appearance adapters share the same architectural form, their parameters are factor-specific and not shared across intrinsic heads. This design allows each adapter to learn distinct attention patterns tailored to its target factor. In practice, we observe that the albedo adapter assigns higher attention weights to geometrically stable regions across views, whereas the diffuse and specular shading adapters attend more strongly to view-dependent highlights and illumination variations. As a result, the adapters effectively decouple material and illumination information before prediction, reducing cross-factor interference and improving both single-view accuracy and multi-view consistency.
@@ 4 Experiments
Table 1: Single-view intrinsic decomposition results on HyperSim.
Method Albedo PSNR $\uparrow$ Albedo MAE $\downarrow$ Albedo SSIM $\uparrow$ Shading PSNR $\downarrow$ Shading SSIM $\uparrow$
IID 15.42 0.061 0.781 - -
RGBX 21.10 0.024 0.742 15.42 0.734
IDT (Ours) 22.85 0.021 0.842 18.32 0.801
Table 2: Single-view depth estimation results.
Method AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}\uparrow$
Depth Anything 0.406 0.391 0.372
VGGT 0.383 0.354 0.412
IDT (Ours) 0.358 0.341 0.433
Table 3: Single-view surface normal estimation results.
Method Mean Angular Error $\downarrow$ $11.25^{\circ}$ $\uparrow$
RGB-X 19.8 58.8
IDT (Ours) 14.1 60.8
Figure 4: Qualitative intrinsic decomposition and geometry prediction results. IDT performs feed-forward inference to predict depth, surface normals, diffuse reflectance (albedo), and diffuse shading from one or multiple views. Compared to existing single-view and multi-view baselines, IDT produces sharper geometry, cleaner intrinsic factors, and substantially improved cross-view consistency.
We evaluate the proposed Intrinsic Decomposition Transformer (IDT) from two complementary perspectives: (i) single-view intrinsic and geometric prediction quality, and (ii) multi-view consistency of the predicted factors. This experimental design explicitly disentangles per-view estimation accuracy from cross-view coherence, which is critical for physically meaningful intrinsic decomposition.
@@ 4.1 Datasets
Hypersim.
Hypersim [Roberts2021Hypersim] is a large-scale synthetic indoor dataset with physically based rendering and full intrinsic ground truth, including depth, surface normals, diffuse reflectance, diffuse illumination, and non-Lambertian residuals. We use Hypersim as the primary benchmark for quantitative evaluation of single-view accuracy and multi-view consistency under controlled supervision.
InteriorVerse.
InteriorVerse [zhu2022learning] contains real-world or photorealistic indoor scenes with complex geometry, materials, and lighting. While full intrinsic ground truth is unavailable, InteriorVerse serves as a challenging testbed for evaluating generalization, visual quality, and cross-view consistency in realistic scenarios.
For both datasets, we construct multi-view samples by grouping $V$ images captured from nearby viewpoints of the same static scene.
@@ 4.2 Evaluation Metrics
Single-view prediction accuracy.
For depth and surface normals, we report standard geometric metrics, including absolute relative error (AbsRel), root mean square error (RMSE), and mean angular error for normals. For intrinsic factors (albedo and shading), we evaluate MAE, PSNR, and SSIM. Shading-related metrics are additionally computed in the logarithmic domain to account for high dynamic range.
@@
To evaluate cross-view consistency, we warp predicted intrinsic factors (albedo, shading, and normals) from each view to a reference view using known camera geometry. We then compute the average $\ell_{1}$ difference between warped predictions. Lower values indicate better multi-view consistency.
@@
We additionally evaluate image reconstruction quality by recomposing the input image using the predicted intrinsic factors according to Eq. (2). PSNR and SSIM are reported on both Hypersim and InteriorVerse.
@@ 4.3 Baselines
We compare IDT against strong baselines tailored to each prediction task.
Depth estimation.
We compare against Depth Anything [depth_anything_v2], VGGT [Wang2025VGGT], which represent state-of-the-art single-view depth estimation methods. All baselines are evaluated in a single-view setting.
Surface normal estimation.
We compare against RGB-X based normal estimation methods [conf/siggraph/0005DGHHLYH24] and intrinsic image decomposition (IID) [Kocsis2023IntrinsicDiffusion] based approaches, which jointly reason about reflectance and shading cues for normal prediction.
Intrinsic image decomposition.
For albedo and shading, we compare against representative intrinsic decomposition methods, including IID models and recent learning-based approaches. All single-view baselines are applied independently to each view.
Multi-view baselines.
To evaluate multi-view consistency, all single-view baselines are extended by independently processing each view, followed by geometric warping. In addition, we include a per-view VGGT baseline that shares the same backbone and training setup as IDT but without joint multi-view attention, allowing us to isolate the effect of joint inference.
4.4 Baselines
We compare IDT against strong baselines tailored to each prediction task.
Depth estimation.
We compare against Depth Anything [depth_anything_v2], VGGT [Wang2025VGGT], which represent state-of-the-art single-view depth estimation methods. All baselines are evaluated in a single-view setting.
Surface normal estimation.
We compare against RGB-X based normal estimation methods [conf/siggraph/0005DGHHLYH24] and intrinsic image decomposition (IID) based approaches, which jointly reason about reflectance and shading cues for normal prediction.
Intrinsic image decomposition.
For albedo and shading, we compare against representative intrinsic decomposition methods, including IID-based models and recent learning-based approaches. All single-view baselines are applied independently to each view.
Multi-view baselines.
To evaluate multi-view consistency, all single-view baselines are extended by independently processing each view, followed by geometric warping. In addition, we include a per-view VGGT baseline that shares the same backbone and training setup as IDT but without joint multi-view attention, allowing us to isolate the effect of joint inference.